At what rate is the volume of the cube changing if the edge is 8cm and is changing at 4cm/sec

1
Expert's answer
2022-04-04T14:26:34-0400

Solution;

Given;

a=8cm

"\frac{da}{dt}=4cm/s"

Since ;

"V=a^3"

"\frac{dV}{dt}=\frac{d}{dt}(a^3)"

"\frac{dV}{dt}=3a^2\frac{da}{dt}"

By direct substitution;

"\frac{dV}{dt}=3(8^2)(4)"

"\frac{dV}{dt}=768cm^3/s"